# Memory-induced current reversal of Brownian motors

### v1

@@ II.2 Quantity of interest
[…]
@@ III Results
In Fig. 3 we present the probability distribution $P(\mathsf{v}(t))$ of $\mathsf{v}(t)$ in the memoryless limit $\tau\to 0$ calculated for $655362^{16}=$ 65536 trajectories and time $t=10^{4}$ for which the asymptotic state of the system has already been reached. Velocity $\mathsf{v}(t)$ is distributed around three values corresponding to average velocities of the particle for three attractors present in the deterministic limit of vanishing thermal noise $D\to 0$ (not depicted). Two $\mathsf{v}_{\pm}=\pm\omega/(2\pi)$ of these values render running solutions in which the particle covers one spatial period during one period of the driving force. The third one $\mathsf{v}_{0}=0$ indicates the locked state in which the particle motion is limited to one potential well and no net transport is observed.

### v2

@@ II.2 Quantity of interest
where the triangular brackets $\langle\cdot\rangle$ indicate averaging over the initial conditions and realizations of thermal noise $\eta(t)$. The former is mandatory for the vanishing and low temperature regimes of the system for which its ergodicity may be broken [88] and therefore the dynamics might depend on the initial conditions.
@@ III Results
In Fig. 3 we present the probability distribution $P(\mathsf{v}(t))$ of $\mathsf{v}(t)$ in the memoryless limit $\tau\to 0$ calculated for $2^{16}=65536$ trajectories and time $t=10^{4}$ for which the asymptotic state of the system has already been reached. Velocity $\mathsf{v}(t)$ is distributed around three values corresponding to average velocities of the particle for three attractors present in the deterministic limit of vanishing thermal noise $D\to 0$ (not depicted). Two $\mathsf{v}_{\pm}=\pm\omega/(2\pi)$ of these values render running solutions in which the particle covers one spatial period during one period of the driving force. The third one $\mathsf{v}_{0}=0$ indicates the locked state in which the particle motion is limited to one potential well and no net transport is observed.
@@ IV Conclusions
Acknowledgments
This work has been supported by the National Science Centre (NCN) under Grant No. 2022/45/B/ST3/02619 (J.S.).